On a metre scale having reading 1cm,2cm,3cm..., masses of 1g,2g,3g,... are kept respectively


if an axis passes through the mid then wat is the moment of inertia of this body

Grade:11

2 Answers

vikas askiitian expert
13 years ago

I = mr2

particles are placed at 1,2,3........& so on

I1 = m1r12 , I2 = m2r22  & so on ........

I = I1 + I2 + I3 ............... n

I = 10-7[(1)(1)2 + 2(2)2 + (3)(32)  ...................n]            ( putting m in grams & r in m)

I =  10-7[13 + 23 + 33  ..........n]

I = [n(n+1)/2 ]2 *10-7 kg-m2

here n is the number of particles ....

Jitender Singh
13 years ago

this will form a series of 100(1 +2^2 +3^2 +.................50^2) soling wil give 0.8585 kg-m^2